Grade upto college level Physical Chemistry

One litre of a mixture of O2 and O3 at NTP was allowed to react with an excess of acidified solution of KI. The iodine liberated required 40 ml of M/10 sodium thiousulphate solution for titration. What is the weight percent of ozone in the mixture? Ultraviolet radiation of wavelength 300 nm can decompose ozone. Assuming that one photon can decompose one ozone molecule, how many photons would have been required for the complete decomposition of ozone in the original mixture?

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
The concerned chemical reaction are :
O3 + 2 KI + H2O → 2KOH + I2 + O­2
I2 + 2Na2S2O3 → Na2S4O6­ + 2Nal
Millimoles of ozone = Millimoles of I2
mM of O3 = mM of I2 = 1/2 * mM of Na2S23
= 1/2 * 40 * 1/10 = 2mM = 0.002 mole
Calculation of total number of moles of O2 and O3
PV = n R T
1 * 1 = n *0.0821 *273 or n = 1/0.0821 *273
Or n = 0.044 mole
∴ Moles of O2 = 0.044 – 0.002 = 0.042
∴ Wt. of O2 = No. of moles * Mol. Wt. = 0.042 * 32 = 1.344 g
Similarly, Wt. of O3 = 0.002 * 48 = 0.096 g
∴ weight % of O3 = 0.096/1.44 * 100 = 6.6%
[Total weight = 1.344 + 0.096 = 1.44 g]
No. of photons or molecules of O3 = 0.096 *6.023 *1023/48 = 1.2 *1021